OpenStudy (anonymous):

Solve. 8x + 4 ≤ –20 A. x ≥ –3 B. x ≥ –2 C. x ≤ –3 D. x ≤ –2

OpenStudy (anonymous):

@DanJS @dtan5457 @jim_thompson5910

OpenStudy (anonymous):

(-20) - 4 = -24 / 8 = -3 and since 8 is not a negative, you keep the sign the same. do you get it?

OpenStudy (anonymous):

so @silverr A or C

OpenStudy (anonymous):

@_alex_urena_

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

C because you keep the same sign, okay ?

OpenStudy (anonymous):

oh silver already got it cx

OpenStudy (anonymous):

thank you :D